\vspace{4mm}

\begin{thebibliography}{99}

\vspace{3mm}


\bibitem{Astashkin1}
{\sc Astashkin S.V.} {\it On the normability of Marcinkiewicz's classes.}  Mathematical Notes,
B. 81, issue 4, April 2007; (in Russian).

\bibitem{Bennet1}
{\sc Bennett C. and Sharpley R.} {\it Interpolation of operators.}  Orlando, Academic Press Inc.,1988.

\bibitem{Fiorenza3}
 {\sc Fiorenza A., and Karadzhov G.E.} {\it Grand and small Lebesgue spaces and
       their analogs.} Consiglio Nationale Delle Ricerche, Instituto per le
      Applicazioni del Calcoto Mauro Picone, Sezione di Napoli, Rapporto tecnico n.
      272/03, (2005).

\bibitem{Iwaniec2}
 {\sc Iwaniec T., P. Koskela P., and Onninen J.} {\it Mapping of finite distortion:
   Monotonicity and Continuity.}  Invent. Math. 144 (2001), 507-531.

\bibitem{Liflyand1}
{\sc Liflyand E., Ostrovsky E., Sirota L.} {\it Structural Properties of Bilateral Grand Lebesgue Spaces.}
 Turk. J. Math.; {\bf 34} (2010), 207-219.

\bibitem{Kozachenko1}
 {\sc Kozatchenko Yu. V., Ostrovsky E.I.} (1985). The Banach Spaces of
      random Variables of subgaussian type. {\it Theory of Probab. and Math.
      Stat.} (in Russian). Kiev, KSU, {\bf 32}, 43-57.

\bibitem{Krein1}
{\sc Krein S.G., Petunin Yu. I., Senenov E.M. } {\it Interpolation of linear operators.}
M;, Nauka,  1978,  (in Russian).

\bibitem{Kufner1}
{\sc Kufner A., John O., Fuchik S. } {\it Function Spaces.}  Noordhoff International Publishing,
Leiden, (1977), Academia; Publishing House of the Czechoslovak Academy of Sciences, Prague.

\bibitem{Kufner2}
{\sc Kufner A.,  Maligranda L, Persson L.-E.} {\it The Hardy Inequality. About its History and Some
     Related Results. } Vydavatelský Servis, Plzen, (2007). 162 pp. ISBN: 978-80-86843-15-5
    MR2351524 (2008j:26001)

\bibitem{Lieb1}
{\sc Lieb E., Loss M.} {\it Analysis.} Providence, Rhode Island, 1997.

\bibitem{Liflyand1}
{\sc Liflyand E., Ostrovsky E., Sirota L.} {\it Structural Properties of Bilateral
Grand Lebesgue Spaces.} Turk. J. Math.; 34 (2010), 207-219.

\bibitem{Okikiolu1}
{\sc Okikiolu G.U.} {\it Aspects of the Theory of Bounded Integral Operators in $ L^p - $  Spaces.}
 Academic Press, London-New York, (1971).

\bibitem{Osekowski1}
{\sc Osekowski A.} {\it Best constants in the weak - type estimates for uncentered maximal operators.}
 Glasgow Mathematical Journal,  {\bf 54,} (2012), pp. 655-663.

\bibitem{Osekowski2}
{\sc Osekowski A.} {\it Sharp $  L^{p, \infty} \rightarrow L^q $  Estimates for the Dyadic-Like Maximal Operators.}
 Journal of Fourier Analysis and Applications,
October 2014, Volume 20, Issue 5, pp. 911-933,  28 May 2014.

\bibitem{Ostrovsky1}
{\sc Ostrovsky E.I.} {\it Exponential Estimations for Random Fields.}
Moscow-Obninsk, OINPE, 1999, ( in Russian).

\bibitem{Ostrovsky2}
{\sc Ostrovsky E.I.} {\it Nikolskii-type inequalities for rearrangement invariant spaces.}
arXiv:0804.2311v1 [math.FA] 15 Apr 2008

\bibitem{Ostrovsky9}
{\sc  Ostrovsky E. and Sirota L.}
{\it A Banach rearrangement norm characterization for tail behavior of measurable functions (random variables) }.
arXiv:1210.1168v1 [math.FA] 3 Oct 2012

\bibitem{Ostrovsky10}
{\sc  Ostrovsky E. and Sirota L.} {\it Central Limit Theorem in H\"older spaces the terms of majorizing measures.}
arXiv:1409.6054v1 [math.PR] 21 Sep 2014

\bibitem{Ostrovsky11}
{\sc  Ostrovsky E. and Sirota L.} {\it Moment Banach spaces: theory and applications.}
HIAT Journal of Science and Engineering, C, Volume 4, Issues 1-2, pp. 233-262, (2007).

\vspace{4mm}

\end{thebibliography}

\documentstyle[12pt]{article}
\newcommand{\const}{\mathop{\rm const}\limits}

\newcommand{\supp}{\mathop{\rm supp}\limits}

\begin{document}

\begin{center}

\vspace{3mm}

{\bf SHARP LYAPUNOV'S INEQUALITY FOR THE }\par

\vspace{4mm}

{\bf MEASURABLE SETS WITH INFINITE MEASURE,} \\

\vspace{4mm}

{\bf with generalization to the Grand Lebesgue spaces.}

\vspace{4mm}

 $ {\bf E.Ostrovsky^a, \ \ L.Sirota^b } $ \\

\vspace{4mm}

$ ^a $ Corresponding Author. Department of Mathematics and computer science, Bar-Ilan University, 84105, Ramat Gan, Israel.\\

E-mail:  \  eugostrovsky@list.ru\\

\vspace{3mm}

$ ^b $  Department of Mathematics and computer science. Bar-Ilan University,
84105, Ramat Gan, Israel.\\

E-mail: \ sirota3@bezeqint.net \\

\vspace{4mm}
                    {\sc Abstract.}\\

 \end{center}

 \vspace{3mm}

  We extend the classical Lyapunov inequality  on the  measurable space with infinite measure and on the
so-called Grand Lebesgue spaces (GLS).\par

 We find also the exact value for correspondent constant. \par

 Possible applications: Functional Analysis (for instance, interpolation of operators), Integral Equations,
Probability Theory and Statistics (tail estimations for random variables) etc.\par

 \vspace{4mm}

{\it Key words and phrases:}  Lebesgue-Riesz and Grand Lebesgue spaces (GLS), measurable set, double ratio
between rearrangement invariant spaces,
fundamental function, upper and lower bounds, rearrangement invariant space and norm, localized GLS norm,
 measurable function (random variable), Lyapunov's inequality. \par

\vspace{3mm}

{\it Mathematics Subject Classification (2000):} primary 60G17; \ secondary
 60E07; 60G70.\\

\vspace{4mm}

\section{Notations. Statement of problem.}

\vspace{4mm}

 Let $ (X = \{x\}, {\bf B}, \mu) $ be measurable space with  non-trivial sigma-finite measure $ \mu. $ We denote as ordinary
for arbitrary measurable numerical (or complex valued) function $ f: X \to R $  and for any measurable set $ A: A \in {\bf B} $
with finite  non-zero measure: $  0 < \mu(A) < \infty $

$$
|f|_{p,A} = \left[  \int_A |f(x)|^p \ d \mu(x) \right]^{1/p},\eqno(1.1)
$$

$$
|f|_{p} = \left[ \int_X |f(x)|^p \ d \mu(x) \right]^{1/p}, \ 1 \le p < \infty,
$$

$$
f \in L_p(A) \ \Leftrightarrow |f|_{p,A} < \infty, \  f \in L_p \ \Leftrightarrow |f|_{p} < \infty.
$$

 Note that $  |f|_{p} = |f|_{p,X}. $ \par

 The classical Lyapunov's inequality asserts  that if  $ \mu(X) = 1, \ 1 \le p \le q,  $  then \\
 $ |f|_p \le |f|_q.  $ Therefore under at the same restrictions on the variables $ p,q $ there holds the following inequality

$$
\frac{|f|_{p,A}}{ \mu^{1/p}(A)} \le  \frac{|f|_{q,A}}{\mu^{1/q}(A) }, \hspace{5mm} 0 < \mu(A) < \infty. \eqno(1.2)
$$

\vspace{4mm}

 {\bf  Our purpose in this short report is to generalize the Lyapunov's inequality (1.2)
 into the so-called Grand Lebesgue Spaces (GLS) instead the classical Lebesgue-Riesz spaces
 constructed over the measurable space with (in general case) infinite measure.   }\par

\vspace{4mm}

 We must  recall here briefly  the definition and some simple properties of the so-called Grand Lebesgue spaces;   more detail
investigation of these spaces see in \cite{Fiorenza3}, \cite{Iwaniec2}, \cite{Kozachenko1}, \cite{Liflyand1}, \cite{Ostrovsky1},
\cite{Ostrovsky2}; see also reference therein.\par

  Recently  appear the so-called Grand Lebesgue Spaces $ GLS = G(\psi) =G\psi =
 G(\psi; a,b), \ a,b = \const, a \ge 1, a < b \le \infty, $ spaces consisting
 on all  the measurable functions $ f: X \to R $ with finite norms

$$
   ||f||G(\psi) \stackrel{def}{=} \sup_{p \in (a,b)} \left[ |f|_p /\psi(p) \right]. \eqno(1.3)
$$

  Here $ \psi(\cdot) $ is some continuous positive on the {\it open} interval
$ (a,b) $ function such that

$$
     \inf_{p \in (a,b)} \psi(p) > 0, \ \psi(p) = \infty, \ p \notin (a,b).
$$

 We will denote
$$
 \supp (\psi) \stackrel{def}{=} (a,b) = \{p: \psi(p) < \infty, \}
$$

The set of all $ \psi $  functions with support $ \supp (\psi)= (a,b) $ will be
denoted by $ \Psi(a,b),  $ and and denote the set of all such a functions as $  \Psi: $

$$
\Psi \stackrel{def}{=} \cup_{1 \le a < b \le \infty} \Psi(a,b).
$$

  The Grand Lebesgue spaces (GLS) are rearrangement invariant, see \cite{Bennet1}, \cite{Krein1}, \cite{Kufner1}, \cite{Lieb1},
and  are used, for example, in the theory of probability \cite{Kozachenko1},  \cite{Ostrovsky1}, \cite{Ostrovsky2}; theory of Partial
Differential Equations \cite{Fiorenza3}, \cite{Iwaniec2};  functional analysis \cite{Fiorenza3}, \cite{Iwaniec2},  \cite{Liflyand1},
 \cite{Ostrovsky2}; theory of Fourier series, theory of martingales, mathematical statistics, theory of approximation etc.\par

\vspace{3mm}

 The so-called {\it fundamental function} $ \phi(\delta) = \phi(G\psi(a,b), \delta), \ \delta \in (0, \infty)  $
  of the $ G\psi(a,b) $ space  may be calculated by the formula

$$
\phi(G\psi(a,b), \delta) = \sup_{p \in (a,b)} \left[ \frac{\delta^{1/p}}{\psi(p)}  \right].
$$

 This notion play a very important role in the theory of interpolation of operators, Fourier series
\cite{Bennet1},   theory of random variables, in particular, theory of Central Limit Theorem,
in Banach spaces \cite{Ostrovsky10} etc. \par
 The detail investigation of fundamental  functions for Grand Lebesgue Spaces with consideration of some examples
see in the articles \cite{Liflyand1}, \cite{Ostrovsky11}. \par

 Let us introduce also the  so-called {\it localized} GLS norm. Indeed,  we define for $ \psi \in \Psi(a,b), \
 A: A \in {\bf B}, \ \mu(A) \in (0, \infty), $ and measurable function $  f: A \to R $

 $$
 ||f||_{\psi,A} := \sup_{p \in (a,b) } \left[ \frac{|f|_{p,A}}{\psi(p)} \right] =
 \sup_{p \in (a,b) } \left[ \frac{[\int_A |f(x)|^p \mu(dx) ]^{1/p}}{\psi(p)} \right]. \eqno(1.4)
 $$

\vspace{4mm}

\section{Main result.}

\vspace{4mm}

 Let $ \psi, \nu $ be two function from the set $  \Psi. $

\vspace{3mm}

{\bf Definition  2.1.} The following functional $ R(G\psi, G\nu) = $

$$
R(\psi, \nu) \stackrel{def}{=} \sup_{ 0 \ne f \in G\psi \cap G\nu}
\sup_{A: 0 < \mu(A) < \infty} \left[ \frac{||f||_{\psi,A}}{\phi(G(\psi), \mu(A))}: \frac{||f||_{\nu,A}}{\phi(G(\nu), \mu(A))} \right] \eqno(2.1)
$$
will named  {\it double ratio}  between the spaces $ G\psi $ and $ G\nu. $ \par

\vspace{3mm}

{\bf Theorem 2.1.} {\it Let $ \psi, \ \nu $ be two function such that $ \psi \in \Psi(a_1,b_1), \ \nu \in \Psi(a_2, b_2). $
Suppose $ b_1 < a_2; $ the opposite case is trivial for us.  Our statement: }

$$
R(G\psi, G\nu) =  1. \eqno(2.2)
$$

\vspace{3mm}

{\bf Proof.} \\

 {\bf First step.} Let $ f \in G\psi \cap G\nu, f \ne 0, \ p \in (a_1,b_1), \ g \in (a_2,b_2), \ 0 < \mu(A) < \infty.  $
We rewrite the Lyapunov's inequality (2.1) as follows:

$$
|f|_{p,A}  \le  \frac{|f|_{q,A}}{\mu^{1/q}(A) } \cdot \mu^{1/p}(A),
$$
or after dividing over $  \psi(p) $

$$
\frac{|f|_{p,A}}{ \psi(p) }  \le  \frac{|f|_{q,A}}{\mu^{1/q}(A) } \cdot \frac{ \mu^{1/p}(A)}{ \psi(p) }. \eqno(2.3)
$$

 We have taking the supremum over $ p, \ p \in (a_1, b_1) $  from both the sides of inequality (2.3)
 using the direct definition of the Grand Lebesgue norm and also the definition of the fundamental function

$$
||f||_{\psi,A} \le \frac{|f|_{q,A}}{\mu^{1/q}(A) } \cdot \phi(G\psi,\mu(A)). \eqno(2.4)
$$

\vspace{3mm}

{\bf Second step.}  Further, we use the simple estimate $ |f|_{q,A} \le ||f||_{\nu,A}\cdot \nu(q), \ q \in (a_2,b_2), $
in the inequality (2.4):

$$
\frac{||f||_{\psi,A}}{ \phi(G\psi,\mu(A)) } \le ||f||_{\nu,A} \cdot \frac{\nu(q)}{\mu^{1/q}(A)}  =
$$

$$
||f||_{\nu,A} \cdot \left[ \frac{\mu^{1/q}(A)}{ \nu(q)  }  \right]^{-1},
$$
and after taking infinum over $  q: $

$$
\frac{||f||_{\psi,A}}{ \phi(G\psi,\mu(A)) } \le ||f||_{\nu,A} \cdot \inf_{q \in (a_2,b_2)} \left[ \frac{\nu(q)}{\mu^{1/q}(A)} \right]  =
$$

$$
||f||_{\nu,A} \cdot \left[ \sup_{q \in (a_2,b_2)} \frac{\mu^{1/q}(A)}{\nu(q)} \right]^{-1} = \frac{||f||_{\nu,A}}{\phi(G\nu,\mu(A))},
$$
hence $  R(G\psi, G\nu) \le 1. $ \\

\vspace{3mm}

{\bf Lower bound.} It is easy to verify that the lower bound for the expression for the functional $ R(G\psi, G\nu) $
is attained, for example, if $  f_0(x) = 1, $ so that $  R(G\psi, G\nu)  \ge 1. $ \par

In detail,

$$
\frac{||f_0||_{\psi,A}}{\phi(G\psi,\mu(A))} =
\frac{\sup_{p \in (a_1,a_2)} \left\{ \left[ \int_A 1 \cdot \mu(dx)  \right]^{1/p}/\psi(p) \right\} }{\phi(G\psi,\mu(A))} =
$$

$$
\frac{\sup_{p \in (a_1,a_2)} \left\{ \left[ \mu(A)\right]^{1/p}/\psi(p) \right\} }{\phi(G\psi, \mu(A))} =
\frac{\phi(G\psi, \mu(A))}{\phi(G\psi, \mu(A))}= 1
$$
and analogously $ ||f_0||_{\nu,A} /\phi(G\nu,A) = 1.$ \par

\vspace{3mm}

 This completes the proof of theorem 2.1.\\

\vspace{4mm}

{\bf Remark 2.1.} The definition (2.1) may be easily extended on  arbitrary pair of rearrangement invariant (r.i.) spaces
$ F_1  $ and $ F_2 $ with norms correspondingly $  ||f||F_1, \ ||f||F_2,  $
constructed over source triple $  (X, {\bf B,} \mu). $ Namely, denote

$$
||f||_{F_1,A} = ||f \cdot I_A||F_1, \hspace{6mm}  ||f||_{F_2,A} = ||f \cdot I_A||F_2,
$$
where as usually $ I_A = I_A(x), \ x \in X $ is indicator function  generated by the measurable set $  A. $ Then by definition of
the double ratio  between the spaces $ F_1 $ and $ F_2 $  is the following functional

$$
R(F_1, F_2) \stackrel{def}{=} \sup_{ 0 \ne f \in F_1 \cap F_2}
\sup_{A: 0 < \mu(A) < \infty} \left[ \frac{||f||_{F_1,A}}{\phi(F_1, \mu(A))}: \frac{||f||_{F_2,A}}{\phi(F_2, \mu(A))} \right].
$$

 Obviously,  $ R(F_1, F_2) \ge 1 $ for any pair of r.i. spaces $ F_1, \ F_2. $ \par

\vspace{3mm}

{\bf Remark 2.2.} If we introduce the {\it discontinuous} function

$$
\psi_r(p) = 1, \ p = r; \psi_r(p) = \infty, \ p \ne r, \ p,r \in (a,b)
$$
and define formally  $ C/\infty = 0, \ C = \const \in R^1, $ then  the norm
in the space $ G(\psi_r) $ coincides with the $ L_r $ norm:

$$
||f||G(\psi_r) = |f|_r.
$$

 Thus, the inequality (1.2) is particular, more exactly, extremal case of the assertion of theorem 2.1.\\

\vspace{4mm}

\section{Concluding remarks.}

 \vspace{4mm}

 Our considerations are very similar to ones in the article \cite{Ostrovsky2}, devoted to the generalization of
Nilol'skii inequality, and may be used perhaps in   the theory of operators acting in the Lorentz spaces, see, e.g.
\cite{Astashkin1}, \cite{Osekowski1}, \cite{Osekowski2}, \cite{Kufner2}, \cite{Okikiolu1}.\\

\end{document}